A right-circular cone is inscribed in a sphere having a fixed radius of 10 in. Express the volume of the cone as a function of its radius.
1
Expert's answer
2020-10-07T16:51:29-0400

The volume of the right circular cone is calculated by the formula.

"V=\frac 1 3\pi r^2H"

where r is the radius and H is the height of the right-circular cone and R=10 sphere radius.

"H=R+\sqrt{\smash[b]{R^2-r^2}}" or "H=R-\sqrt{\smash[b]{R^2-r^2}}"

Hence the function of the volume of the cone from its base is given as

"V=\frac 1 3\pi r^2H=\frac 1 3\pi r^2×(R±\sqrt{\smash[b]{R^2-r^2}})"


"V=\frac 1 3\pi r^2×(10±\sqrt{\smash[b]{100-r^2}})" Cubic Inches